OpenStudy (jadonbb):

Can you guys explain how to solve this (x/x-2)+(x-1/x+1)=-1

OpenStudy (jadonbb):

\[(x \div x-2) + (x-1\div x+1)=-1\]

OpenStudy (p0sitr0n):

so um first put them all on a common denominator: \(\frac{2x^2-2x+2}{(x-2)(x+1)}=-1 \) or \( 2x^2-2x+2 = (2-x)(x+1)\). Now solve it.

OpenStudy (mathstudent55):

Although you wrote (x/x-2)+(x-1/x+1)=-1 which means \(\dfrac{x}{x}-2+x-\dfrac{1}{x}+1=-1 \) you really meant this \(\dfrac{x}{x-2}+\dfrac{x-1}{x+1}=-1 \) right?

OpenStudy (jadonbb):

yes @mathstudent55

OpenStudy (mathstudent55):

If the problem is really the last one above, then there are several ways of solving. Here is one of them. Add the fractions on the left side by using the LCD. The LCD of x - 2 and x + 1 is (x - 2)(x + 1).

OpenStudy (jadonbb):

@mathstudent55 I actually solved it earlier today x=0,1 is i remember correctly

OpenStudy (mathstudent55):

We need to multiply the left fraction by \(\dfrac{x + 1}{x + 1}\) and the right fraction by \(\dfrac{x - 2}{x - 2} \) to have a common denominator. \(\dfrac{x}{x-2} \times \dfrac{x + 1}{x + 1} +\dfrac{x-1}{x+1} \times \dfrac{x - 2}{x - 2} =-1\) \(\dfrac{x(x + 1)}{(x+1)(x - 2)} +\dfrac{(x-1)(x - 2)}{(x+1)(x - 2)} =-1\) \(\dfrac{x^2 + x}{(x+1)(x - 2)} +\dfrac{x^2 -2x - x + 2}{(x + 1)(x - 2)} =-1\) \(\dfrac{2x^2 -2x + 2}{(x+1)(x - 2)} =-1\) \(2x^2 -2x + 2 =-1(x+1)(x - 2)\) \(2x^2 -2x + 2 = -(x^2 -2x + x - 2)\) \(2x^2 -2x + 2 = -x^2 + x + 2\) \(3x^2 - 3x = 0\) \(3x(x - 1) = 0\) \(x = 0\) or \(x - 1 = 0\) \(x = 0\) or \(x = 1\)

OpenStudy (mathstudent55):

Now we check the original equation. It has denominators x - 2 and x + 1. That means the solution cannot be x = 2 or x = -1. Since our solutions are x = 0 and x = 1, these solutions are acceptable, and they are the correct solution to the equation.